OpenStudy (anonymous):

4(x + 6) = 6(x − 2) I did: 4(x)+4(6)=6(x)-6(2) 4x+24=6x-12 10x+24-12 10x+12 Then I got stuck.

OpenStudy (lgbasallote):

why did you lose an equal sign?

OpenStudy (anonymous):

oh! i see what I did. 4x+24=6x-12 10x+24=-12 10x=-36

OpenStudy (asnaseer):

your step after this one: 4x+24=6x-12 is wrong. you need to subtract 6x from both sides. then subtract 24 from both sides.

OpenStudy (lgbasallote):

yup that was the second mistake

OpenStudy (anonymous):

so i subtract 6x from 4x?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Why is it that i subtract 6x?

OpenStudy (asnaseer):

think of it like this: 4x + 24 = 6x - 12 4x + 24 - 6x = 6x - 12 -6x <-- this gets rid of the 6x from the right-hand-side

OpenStudy (asnaseer):

the idea is to keep the equation "balanced". so whatever you do to one side, you also need to do to the other side.

OpenStudy (anonymous):

I got x= -18 and thank you for explaining that!

OpenStudy (asnaseer):

yw :) however, I don't believe you have the right answer.

OpenStudy (asnaseer):

what did you get after subtracting 6x from both sides?

OpenStudy (anonymous):

i got: 2x+24=-12 -24 -24 2x=-36 ------- 2x 2x x= -18

OpenStudy (asnaseer):

4x - 6x does not equal 2x ---> you are missing a minus sign

OpenStudy (anonymous):

sorry typo. i did do -2x on paper.

OpenStudy (asnaseer):

so you should have ended up with: -2x + 24 = -12

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

now subtract 24 from both sides - what do you end up with?

OpenStudy (asnaseer):

are you stuck?

OpenStudy (anonymous):

-2x=-36 -------- -2x -2x x=18 The site locked up.

OpenStudy (asnaseer):

thats the correct answer now - well done! ::)

OpenStudy (anonymous):

yay!! thank you again!

OpenStudy (asnaseer):

yw :)